In each part, write the set in the form {x|P(x)}, where P(x) is a property that describes the
elements of the set.
a. {2, 4, 6, 8, 10}
b. {a, e, o, i, u}
c. {1, 4, 9, 16, 25, 36}
d. {-2, -1, 0, 1, 2}

Answer by oscargut(2103)

a. {2, 4, 6, 8, 10} = {x / x is even 2<=x<=10}
b. {a, e, o, i, u} = {x/ x is a vowel}
c. {1, 4, 9, 16, 25, 36} = {x / x =n^2 , 1<=x<=6}
d. {-2, -1, 0, 1, 2} = {x / x in Z / -2<=x<=2}
You can ask me more at: mthman@gmail.com
Thanks